Differences of squares pattern to factor the following:
(x+6)^2 - (x+9)^2

Answer by jake_6233(74)

(x+6)^2 - (x+9)^2
(a+b)(a-b) = a^2 - b^2 where a= (x+6) b=(x+9)
((x+6) + (x+9))-((x+6) - (x+9))